GNU Automake

For version 1.0, 25 May 1996 Copyright \begin{matrix}c \\ ◯ \end{matrix} 1995 Free Software Foundation, Inc.

This is the first edition of the GNU Automake documentation, and is consistent with GNU Automake 1.0.

Published by the Free Software Foundation 675 Massachusetts Avenue, Cambridge, MA 02139 USA

Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies.

Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one.

Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that this permission notice may be stated in a translation approved by the Free Software Foundation.

1 Introduction

Automake is a tool for automatically generating ‘Makefile.in’s from files called ‘Makefile.am’. The ‘Makefile.am’ is basically a series of make macro definitions (with rules being thrown in occasionally). The generated ‘Makefile.in’s are compliant with the GNU Makefile standards.

The GNU Makefile Standards Document (see section “Makefile Conventions” in The GNU Coding Standards) is long, complicated, and subject to change. The goal of Automake is to remove the burden of Makefile maintenance from the back of the individual GNU maintainer (and put it on the back of the Automake maintainer).

The typical Automake input files is simply a series of macro definitions. Each such file is processed to create a ‘Makefile.in’. There should generally be one ‘Makefile.am’ per directory of a project.

Automake does constrain a project in certain ways; for instance it assumes that the project uses Autoconf (see section “The Autoconf Manual” in The Autoconf Manual), and enforces certain restrictions on the ‘configure.in’ contents.

Automake requires perl in order to generate the ‘Makefile.in’s. However, the distributions created by Automake are fully GNU standards-compliant, and do not require perl in order to be built.

Mail suggestions and bug reports for Automake to tromey@cygnus.com.

2 Creating a ‘Makefile.in’

To create all the ‘Makefile.in’s for a package, run the automake program in the top level directory, with no arguments. automake will automatically find each appropriate ‘Makefile.am’ (by scanning ‘configure.in’; see Chapter 4 [configure], page 6) and generate the corresponding

‘Makefile.in’.

You can optionally give automake an argument; ‘.am’ is appended to the argument and the result is used as the name of the input file. This feature is generally only used to automatically rebuild an out-of-date ‘Makefile.in’. Note that automake must always be run from the topmost directory of a project, even if being used to regenerate the ‘Makefile.in’ in some subdirectory.

This is necessary because automake must scan ‘configure.in’, and because automake uses the knowledge that a ‘Makefile.in’ is in a subdirectory to change its behavior in some cases.

automake accepts the following options:

--amdir=dir

Look for Automake data files in directory dir instead of in the installation directory. This is typically used for debugging.

--foreign
An alias for ‘--strictness=foreign’.
--gnits An alias for ‘--strictness=gnits’.
--gnu  An alias for ‘--strictness=gnu’.

--help Print a summary of the command line options and exit.

--include-deps

Include all automatically generated dependency information (see Section 6.4 [Dependencies], page 12) in the generated ‘Makefile.in’. This is generally done when making a distribution; see Chapter 11 [Dist], page 16.

--add-missing

Automake requires certain common files to exist in certain situations; for instance

‘config.guess’ is required if ‘configure.in’ runs AC_CANONICAL_HOST. Automake is

distributed with several of these files; this option will cause the missing ones to be automatically added to the package, whenever possible.

--output-dir=dir

Put the generated ‘Makefile.in’ in the directory dir. Ordinarily each ‘Makefile.in’ is created in the directory of the corresponding ‘Makefile.am’. This option is used when making distributions.

--strictness=level

Set the global strictness to level; this can be overridden in each ‘Makefile.am’ if required. See Chapter 3 [Generalities], page 3 for more information.

--verbose

Cause Automake to print information about which files are being read or created.

--version

Print the version number of Automake and exit.

3 General ideas

3.1 Depth

automake supports three kinds of directory hierarchy: “flat”, “shallow”, and “deep”.

A flat package is one in which all the files are in a single directory. The ‘Makefile.am’ for such a package by definition lacks a SUBDIRS macro. An example of such a package is termutils.

A deep package is one in which all the source lies in subdirectories; the top level directory contains mainly configuration information. GNU cpio is a good example of such a package, as is GNU tar. The top level ‘Makefile.am’ for a deep package will contain a SUBDIRS macro, but no other macros to define objects which are built.

A shallow package is one in which the primary source resides in the top-level directory, while various parts (typically libraries) reside in subdirectories. automake is one such package (as is GNU make, which does not currently use automake).

3.2 Strictness

While Automake is intended to be used by maintainers of GNU packages, it does make some effort to accomodate those who wish to use it, but do not want to use all the GNU conventions.

To this end, Automake supports three levels of strictness – the strictness indicating how stringently Automake should check standards conformance.

The valid strictness levels are:

‘foreign’ Automake will check for only those things which are absolutely required for proper operations. For instance, whereas GNU standards dictate the existence of a ‘NEWS’ file, it will not be required in this mode. The name comes from the fact that Automake is intended to be used for GNU programs; these relaxed rules are not the standard mode of operation.

‘gnu’ Automake will check – as much as possible – for compliance to the GNU standards for packages. This is the default.

‘gnits’ Automake will check for compliance to the as-yet-unwritten GNITS standards. These are based on the GNU standards, but are even more detailed. Unless you are a GNITS standards contributor, it is recommended that you avoid this option until such time as the GNITS standard is actually published.

3.3 The Uniform Naming Scheme

Automake variables generally follow a uniform naming scheme that makes it easy to decide how programs (and other derived objects) are built, and how they are installed. This scheme also supports configure time determination of what should be built.

At make time, certain variables are used to determine which objects are to be built. These variables are called primary variables. For instance, the primary variable PROGRAMS holds a list of programs which are to be compiled and linked.

A different set of variables is used to decide where the built objects should be installed. These variables are named after the primary variables, but have a prefix indicating which standard directory should be used as the installation directory. The standard directory names are given in the GNU standards (see section “Directory Variables” in The GNU Coding Standards). automake extends this list with pkglibdir, pkgincludedir, and pkgdatadir; these are the same as the non-‘pkg’ versions, but with ‘@PACKAGE@’ appended.

For each primary, there is one additional variable named by prepending ‘EXTRA_’ to the primary name. This variable is used to list objects which may or may not be built, depending on what configure decides. This variable is required because Automake must know the entire list of objects to be built in order to generate a ‘Makefile.in’ that will work in all cases.

For instance, cpio decides at configure time which programs are built. Some of the programs are installed in bindir, and some are installed in sbindir:

     EXTRA_PROGRAMS = mt rmt
     bin_PROGRAMS = cpio pax
     sbin_PROGRAMS = @PROGRAMS@

Defining a primary variable is an error.

Note that the common ‘dir’ suffix is left off when constructing the variable names; thus one

writes ‘bin_PROGRAMS’ and not ‘bindir_PROGRAMS’.

Not every sort of object can be installed in every directory. Automake will flag those attempts it finds in error. Automake will also diagnose obvious misspellings in directory names.

Sometimes the standard directories – even as augmented by Automake – are not enough. In particular it is sometimes useful, for clarity, to install objects in a subdirectory of some predefined directory. To this end, Automake allows you to extend the list of possible installation directories. A given prefix (eg ‘zar’) is valid if a variable of the same name with ‘dir’ appended is defined (eg

‘zardir’).

For instance, until HTML support is part of Automake, you could use this to install raw HTML documentation:

     htmldir = $(prefix)/html
     html_DATA = automake.html

The special prefix ‘noinst’ indicates that the objects in question should not be installed at all.

The special prefix ‘check’ indicates that the objects in question should not be built until the make check command is run.

Possible primary names are ‘PROGRAMS’, ‘LIBRARIES’, ‘SCRIPTS’, ‘DATA’, ‘HEADERS’, ‘MANS’, and

‘TEXINFOS’.

3.4 General Operation

Automake essentially works by reading a ‘Makefile.am’ and generating a ‘Makefile.in’. The macro definitions and targets in the ‘Makefile.am’ are copied into the generated file.

Automake tries to group comments with adjoining targets (or variable definitions) in an intelligent way.

A target defined in ‘Makefile.am’ generally overrides any such target of a similar name that would be automatically generated by automake. Although this is a supported feature, it is generally best to avoid making use of it, as sometimes the generated rules are very particular.

Automake also allows a form of comment which is not copied into the output; all lines beginning with ‘##’ are completely ignored by Automake.

It is customary to make the first line of ‘Makefile.am’ read:

     ## Process this file with automake to produce Makefile.in

4 Scanning ‘configure.in’

Automake requires certain variables to be defined and certain macros to be used in the package

‘configure.in’.

One such requirement is that ‘configure.in’ must define the variables PACKAGE and VERSION with AC_SUBST. PACKAGE should be the name of the package as it appears when bundled for distribution. For instance, Automake definees PACKAGE to be ‘automake’. VERSION should be the version number of the release that is being developed. We recommend that you make ‘configure.in’ the only place in your package where the version number is defined; this makes releases simpler.

Automake requires the use of the macro AC_ARG_PROGRAM if a program or script is installed.

If your package is not a flat distribution, Automake requires the use of AC_PROG_MAKE_SET.

Automake will also recognize the use of certain macros and tailor the generated ‘Makefile.in’ appropriately. Currently recognized macros and their effects are:

AC_CONFIG_HEADER

Automake will generate rules to automatically regenerate the config header. If you do use this macro, you must create the file ‘stamp-h.in’. It can be empty. Also, the

          AC_OUTPUT command in ‘configure.in’ must create ‘stamp-h’, eg:
               AC_OUTPUT(Makefile, [test -z "$CONFIG_HEADERS" || echo timestamp > stamp-h])
AC_CONFIG_AUX_DIR

Automake will look for various helper scripts, such as ‘mkinstalldirs’, in the directory named in this macro invocation. If not seen, the scripts are looked for in their “standard” locations (either the top source directory, or in the source directory corresponding to the current ‘Makefile.am’, whichever is appropriate).

AC_OUTPUT

Automake uses this to determine which files to create.

AC_PATH_XTRA

Automake will insert definitions for the variables defined by AC_PATH_XTRA into each ‘Makefile.in’ that builds a C program or library.

AC_CANONICAL_HOST
AC_CANONICAL_SYSTEM
AC_CHECK_TOOL

Automake will ensure that ‘config.guess’ and ‘config.sub’ exist.

AC_FUNC_ALLOCA
AC_FUNC_GETLOADAVG
AC_FUNC_MEMCMP
AC_STRUCT_ST_BLOCKS
fp_FUNC_FNMATCH
AC_FUNC_FNMATCH
AC_REPLACE_FUNCS
AC_REPLACE_GNU_GETOPT

Automake will ensure that the appropriate source files are part of the distribution, and will ensure that the appropriate dependencies are generated for these objects. See Section 6.2 [A Library], page 11 for more information. Automake will also detect statements which put ‘.o’ files into LIBOBJS, and will treat these additional files in a similar way.

AC_PROG_RANLIB

This is required if any libraries are built in the package.

AC_PROG_INSTALL
fp_PROG_INSTALL

fp_PROG_INSTALL is required if any scripts (see Section 7.1 [Scripts], page 12) are installed by the package. Otherwise, AC_PROG_INSTALL is required.

gm_PROG_LIBTOOL
AC_PROG_LIBTOOL

Automake will turn on processing for libtool (see section “The Libtool Manual” in The Libtool Manual). This work is still preliminary.

ALL_LINGUAS

If Automake sees that this variable is set in ‘configure.in’, it will check the ‘po’ directory to ensure that all the named ‘.po’ files exist, and that all the ‘.po’ files that exist are named.

fp_C_PROTOTYPES

This is required when using automatic de-ANSI-fication, see Section 6.3 [ANSI], page 11.

ud_GNU_GETTEXT

This macro is required for packages which use GNU gettext (FIXME xref). It is distributed with gettext. Automake uses this macro to ensure that the package meets some of gettext’s requirements.

jm_MAINTAINER_MODE

This macro adds a ‘--enable-maintainer-mode’ option to configure. If this is used, automake will cause “maintainer-only” rules to be turned off by default in the generated

          ‘Makefile.in’s.

The ‘fp_’ macros are from Francois Pinard and the ‘jm_’ macro is from Jim Meyering. Both sets are included with Automake. automake will tell where the macros can be found if they are missing.

5 The top-level ‘Makefile.am’

In non-flat packages, the top level ‘Makefile.am’ must tell Automake which subdirectories are to be built. This is done via the SUBDIRS variable.

The SUBDIRS macro holds a list of subdirectories in which building of various sorts can occur. Many targets (eg all) in the generated ‘Makefile’ will run both locally and in all specified subdirectories. Note that the directories listed in SUBDIRS are not required to contain ‘Makefile.am’s; only ‘Makefile’s (after configuration). This allows inclusion of libraries from packages which do not use Automake (such as gettext).

In a deep package, the top-level ‘Makefile.am’ is often very short. For instance, here is the ‘Makefile.am’ from the textutils distribution:

     SUBDIRS = lib src doc man
     EXTRA_DIST = @README_ALPHA@

SUBDIRS can contain configure substitutions (eg ‘@DIRS@’); Automake itself does not actually examine the contents of this variable.

   If SUBDIRS is defined, then your ‘configure.in’ must include AC_PROG_MAKE_SET.

6 Building Programs and Libraries

A large part of Automake’s functionality is dedicated to making it easy to build C programs and libraries.

6.1 Building a program

In a directory containing source that gets built into a program (as opposed to a library), the ‘PROGRAMS’ primary is used. Programs can be installed in bindir, sbindir, libexecdir, pkglibdir, or not at all.

For instance:

     bin_PROGRAMS = hello

In this simple case, the resulting ‘Makefile.in’ will contain code to generate a program named hello. The variable hello_SOURCES is used to specify which source files get built into an executable:

     hello_SOURCES = hello.c

This causes ‘hello.c’ to be compiled into ‘hello.o’, and then linked to produce ‘hello’.

If ‘prog_SOURCES’ is needed, but not specified, then it defaults to the single file ‘prog.c’. Id est in the example above, the definition of hello_SOURCES is actually redundant.

Multiple programs can be built in a single directory. Multiple programs can share a single source file. The source file must be listed in each ‘_SOURCES’ definition.

Header files listed in a ‘_SOURCES’ definition will be ignored. Lex (‘.l’) and yacc (‘.y’) files can also be listed; support for these should work but is still preliminary.

Sometimes it is useful to determine the programs that are to be built at configure time. For instance, GNU cpio only builts mt and rmt under special circumstances.

In this case, you must notify Automake of all the programs that can possibly be built, but at the same time cause the generated ‘Makefile.in’ to use the programs specified by configure. This is done by having configure substitute values into each ‘_PROGRAMS’ definition, while listing all optionally built programs in EXTRA_PROGRAMS.

If you need to link against libraries that are not found by configure, you can use LDADD to do so. This variable actually can be used to add any options to the linker command line.

Sometimes, multiple programs are built in one directory but do not share the same link-time requirements. In this case, you can use the ‘prog_LDADD’ variable (where PROG is the name of the program as it appears in some ‘_PROGRAMS’ variable, and usually written in lowercase) to override the global LDADD. (If this variable exists for a given program, then that program is not linked using

LDADD.)

For instance, in GNU cpio, pax, cpio, and mt are linked against the library ‘libcpio.a’. However, rmt is built in the same directory, and has no such link requirement. Also, mt and rmt are only built on certain architectures. Here is what cpio’s ‘src/Makefile.am’ looks like (abridged):

     bin_PROGRAMS = cpio pax @MT@
     libexec_PROGRAMS = @RMT@
     EXTRA_PROGRAMS = mt rmt
     LDADD = ../lib/libcpio.a @INTLLIBS@
     rmt_LDADD =
     cpio_SOURCES = . . .
     pax_SOURCES = . . .
     mt_SOURCES = . . .
     rmt_SOURCES = . . .

It is also occasionally useful to have a program depend on some other target which is not actually part of that program. This can be done using the ‘prog_DEPENDENCIES’ variable. Each program depends on the contents of such a variable, but no further interpretation is done.

Since program names are rewritten into Makefile macro names, program names must follow Makefile macro syntax. Sometimes it is useful to have a program whose name does not follow such rules. In these cases, Automake canonicalizes the program name. All characters in the name except for letters, numbers, and the underscore are turned into underscores when making macro references. Eg, if your program is named sniff-glue, you would use sniff_glue_SOURCES, not

sniff-glue_SOURCES.

6.2 Building a library

Building a library is much like building a program. In this case, the name of the primary is ‘LIBRARIES’. Libraries can be installed in libdir or pkglibdir.

Each ‘_LIBRARIES’ variable is a list of the base names of libraries to be built. For instance to create a library named ‘libcpio.a’, but not install it, you would write:

     noinst_LIBRARIES = cpio

The sources that go into a library are determined exactly as they are for programs, via the ‘_SOURCES’ variables. Note that programs and libraries share a namespace, so one cannot have a program (‘lob’) and a library (‘liblob.a’) with the same name in one directory.

Extra objects can be added to a library using the ‘library_LIBADD’ variable. This should be used for objects determined by configure. Again from cpio:

     cpio_LIBADD = @LIBOBJS@ @ALLOCA@

Note that Automake explicitly recognizes the use of @LIBOBJS@ and @ALLOCA@ in the above example, and uses this information, plus the list of LIBOBJS files derived from ‘configure.in’ to automatically include the appropriate source files in the distribution (see Chapter 11 [Dist], page 16). These source files are also automatically handled in the dependency-tracking scheme, see See Section 6.4 [Dependencies], page 12.

6.3 Automatic de-ANSI-fication

Although the GNU standards prohibit it, some GNU programs are written in ANSI C; see FIXME. This is possible because each source file can be “de-ANSI-fied” before the actual compilation takes place.

If the ‘Makefile.am’ variable AUTOMAKE_OPTIONS (Chapter 13 [Options], page 17) contains the option ansi2knr then code to handle de-ANSI-fication is inserted into the generated ‘Makefile.in’.

This causes each source file to be treated as ANSI C. If an ANSI C compiler is available, it is used.

This support requires the source files ‘ansi2knr.c’ and ‘ansi2knr.1’ to be in the same directory as the ANSI C source; these files are distributed with Automake. Also, the package ‘configure.in’ must call the macro fp_C_PROTOTYPES.

6.4 Automatic dependency tracking

As a developer it is often painful to continually update the ‘Makefile.in’ whenever the includefile dependencies change in a project. automake supplies a way to automatically track dependency changes, and distribute the dependencies in the generated ‘Makefile.in’.

Currently this support requires the use of GNU make and gcc. It might become possible in the future to supply a different dependency generating program, if there is enough demand.

This mode is enabled by default if any C program or library is defined in the current directory.

When you decide to make a distribution, the dist target will re-run automake with the ‘--include-deps’ option. This causes the previously generated dependencies to be inserted into the generated ‘Makefile.in’, and thus into the distribution. ‘--include-deps’ also turns off inclusion of the dependency generation code.

This mode can be suppressed by putting no-dependencies in the variable AUTOMAKE_OPTIONS.

7 Other Derived Objects

Automake can handle derived objects which are not C programs. Sometimes the support for actually building such objects must be explicitly supplied, but Automake will still automatically handle installation and distribution.

7.1 Executable Scripts

It is possible to define and install programs which are scripts. Such programs are listed using the ‘SCRIPTS’ primary name. automake doesn’t define any dependencies for scripts; the ‘Makefile.am’ should include the appropriate rules.

automake does not assume that scripts are derived objects; such objects are must be deleted by hand; see Chapter 10 [Clean], page 16 for more information.

automake itself is a script that is generated at configure time from ‘automake.in’. Here is how this is handled:

  bin_SCRIPTS = automake

Since automake appears in the AC_OUTPUT macro, dependencies for it are automatically generated.

Script objects can be installed in bindir, sbindir, libexecdir, or pkgdatadir.

7.2 Header files

Header files are specified by the ‘HEADERS’ family of variables. Generally header files are not installed, so the noinst_HEADERS variable will be the most used.

All header files must be listed somewhere; missing ones will not appear in the distribution. Often it is most convenient to list uninstalled headers with the rest of the sources for a program. See Section 6.1 [A Program], page 9.

Headers can be installed in includedir, oldincludedir, or pkgincludedir.

7.3 Architecture-independent data files

Automake supports the installation of miscellaneous data files using the ‘DATA’ family of variables.

Such data can be installed in the directories datadir, sysconfdir, sharedstatedir, localstatedir,
or pkgdatadir.

All such data files are included in the distribution. Here is how autoconf installs its auxiliary data files:

     pkgdata_DATA = clean-kr.am clean.am compile-kr.am compile-vars.am \
     compile.am data.am depend.am dist-subd-top.am dist-subd-vars.am \
     dist-subd.am dist-vars.am dist.am footer.am header-vars.am header.am \
     libscripts.am libprograms.am libraries-vars.am libraries.am library.am \
     mans-vars.am mans.am packagedata.am program.am programs.am remake-hdr.am \
     remake-subd.am remake.am scripts.am subdirs.am tags.am tags-subd.am \
     texinfos-vars.am texinfos.am hack-make.sed nl-remove.sed

7.4 Built sources

Occasionally a file which would otherwise be called “source” (eg a C ‘.h’ file) is actually derived from some other file. Such files should be listed in the BUILT_SOURCES variable.

Files listed in BUILT_SOURCES are built before any automatic dependency tracking is done. Built sources are included in a distribution.

8 Building documentation

Currently Automake provides support for Texinfo and man pages.

8.1 Texinfo

If the current directory contains Texinfo source, you must declare it with the ‘TEXINFOS’ primary. Generally Texinfo files are converted into info, and thus the info_TEXINFOS macro is most commonly used here. Note that any Texinfo source file must end in the ‘.texi’ extension (‘.texinfo’ won’t work).

If the ‘.texi’ file @includes ‘version.texi’, then that file will be automatically generated. ‘version.texi’ defines three Texinfo macros you can reference: EDITION, VERSION, and UPDATED. The first two hold the version number of your package (but are kept separate for clarity); the last is the date the primary file was last modified. The ‘version.texi’ support requires the mdate-sh program; this program is supplied with Automake.

Sometimes an info file actually depends on more than one ‘.texi’ file. For instance, in

the xdvik distribution, ‘kpathsea.texi’ includes the files ‘install.texi’, ‘copying.texi’, and

‘freedom.texi’. You can tell Automake about these dependencies using the ‘texi_TEXINFOS’ variable. Here is how xdvik could do it:

     info_TEXINFOS = kpathsea.texi
     kpathsea_TEXINFOS = install.texi copying.texi freedom.texi

Automake will warn if a directory containing Texinfo source does not also contain the file ‘texinfo.tex’. This file is supplied with Automake.

Automake generates an install-info target; some people apparently use this.

8.2 Man pages

A package can also include man pages. (Though see the GNU standards on this matter, section “Man Pages” in The GNU Coding Standards.) Man pages are declared using the ‘MANS’ primary. Generally the man_MANS macro is used. Man pages are automatically installed in the correct subdirectory of mandir, based on the file extension.

By default, man pages are installed by ‘make install’. However, since the GNU project does not require man pages, many maintainers do not expend effort to keep the man pages up to date. In these cases, the no-installman option will prevent the man pages from being installed by default. The user can still explicitly install them via ‘make install-man’.

Here is how the documentation is handled in GNU cpio (which includes both Texinfo documentation and man pages):

     info_TEXINFOS = cpio.texi
     man_MANS = cpio.1 mt.1

Texinfo source, info pages and man pages are all considered to be source for the purposes of making a distribution.

9 What Gets Installed

Naturally, Automake handles the details of actually installing your program once it has been built. All PROGRAMS, SCRIPTS, LIBRARIES, DATA and HEADERS are automatically installed in the appropriate places.

Automake also handles installing any specified info and man pages.

Automake generates separate install-data and install-exec targets, in case the installer is installing on multiple machines which share directory structure – these targets allow the machineindependent parts to be installed only once. The install target depends on both of these targets.

Automake also generates an uninstall target, and an installdirs target.

It is possible to extend this mechanism by defining an install-exec-local or install-data-local target. If these targets exist, they will be run at ‘make install’ time.

10 What Gets Cleaned

The GNU Makefile Standards specify a number of different clean rules. Generally the files that can cleaned are determined automatically by Automake. Of course, Automake also recognizes some variables that can be defined to specify additional files to clean. These variables are

MOSTLYCLEANFILES, CLEANFILES, DISTCLEANFILES, and MAINTAINERCLEANFILES.

In Automake, the automake program is not automatically removed, because it is an executable script. So this code in ‘Makefile.am’ causes it to be removed by ‘make clean’:

     CLEANFILES = automake

11 What Goes in a Distribution

The dist target in the generated ‘Makefile.in’ can be used to generate a gzip’d tar file for distribution. The tar file is named based on the PACKAGE and VERSION variables.

For the most part, the files to distribute are automatically found by Automake: all source files are automatically included in a distribution, as are all ‘Makefile.am’s and ‘Makefile.in’s. Automake also has a built-in list of commonly used files which, if present in the current directory, are automatically included. This list is printed by ‘automake --help’. Also, files which are read by configure (ie, the source files corresponding to the files specified in the AC_OUTPUT invocation) are automatically distributed.

Still, sometimes there are files which must be distributed, but which are not covered in the automatic rules. These files should be listed in the EXTRA_DIST variable.

Occasionally it is useful to be able to change the distribution before it is packaged up. If the dist-hook target exists, it is run after the distribution directory is filled, but before the actual tar (or shar) file is created. One way to use this is for distributing file in subdirectories for which a new ‘Makefile.am’ is overkill:

     dist-hook:
             mkdir $(distdir)/random
             cp -p random/a1 random/a2 $(distdir)/random

Automake also generates a distcheck target which can be help to ensure that a given distribution will actually work. distcheck makes a distribution, and then tries to do a VPATH build.

12 Support for test suites

Automake supports a two forms of test suite.

If the variable TESTS is defined, its value is taken to be a list of programs to run in order to do the testing. The programs can either be derived objects or source objects; the generated rule will look both in srcdir and ‘.’. The number of failures will be printed at the end of the run.

If ‘dejagnu’ appears in AUTOMAKE_OPTIONS, then the a dejagnu-based test suite is assumed. The value of the variable DEJATOOL is passed as the --tool argument to runtest; it defaults to the name of the package. The variables EXPECT, RUNTEST and RUNTESTFLAGS can also be overridden to provide project-specific values. For instance, you will need to do this if you are testing a compiler toolchain, because the default values do not take into account host and target names.

In either case, the testing is done via ‘make check’.

13 Changing Automake’s Behavior

Various features of Automake can be controlled by options in the ‘Makefile.am’. Such options are listed in a special variable named AUTOMAKE_OPTIONS. Currently understood options are:

gnits
gnu

foreign The same as the corresponding ‘--strictness’ option.

no-installman

The generated ‘Makefile.in’ will not cause man pages to be installed by default. However, an install-man target will still be available for optional installation.

ansi2knr Turn on automatic de-ANSI-fication.

dist-shar

Generate a dist-shar target as well as the ordinary dist target.

dist-zip Generate a dist-zip target as well as the ordinary dist target.

no-dependencies

This is similar to using ‘--include-deps’ on the command line, but is useful for those situations where you don’t have the necessary bits to make automatic dependency tracking work See Section 6.4 [Dependencies], page 12. In this case the effect is to effectively disable automatic dependency tracking.

version A version number (eg ‘0.30’) can be specified. If Automake is not newer than the version specified, creation of the ‘Makefile.in’ will be suppressed.

Unrecognized options are diagnosed by automake.

14 Miscellaneous Rules

There are a few rules and variables that didn’t fit anywhere else.

14.1 Interfacing to etags

automake will generate rules to generate ‘TAGS’ files for use with GNU Emacs under some circumstances.

If any C source code or headers are present, then a tags target will be generated for the directory.

At the topmost directory of a multi-directory package, a tags target file will be generated which, when run, will generate a ‘TAGS’ file that includes by reference all ‘TAGS’ files from subdirectories.

Also, if the variable ETAGS_ARGS is defined, a tags target will be generated. This variable is intended for use in directories which contain taggable source that etags does not understand.

Here is how Automake generates tags for its source, and for nodes in its Texinfo file:

     ETAGS_ARGS = automake.in --lang=none \
      --regex='/^@node[ \t]+\([^,]+\)/\1/' automake.texi

Automake will also generate an ID target which will run mkid on the source. This is only supported on a directory-by-directory basis.

14.2 Handling new file extensions

It is sometimes useful to introduce a new implicit rule to handle a file type that Automake does not know about. If this is done, you must notify GNU Make of the new suffixes. This can be done by putting a list of new suffixes in the SUFFIXES variable.

14.3 Built sources

FIXME write this

15 When Automake Isn’t Enough

Sometimes automake isn’t enough. Then you just lose.

Actually, automakes implicit copying semantics means that many problems can be worked around by simply adding some make targets and rules to ‘Makefile.in’. automake will ignore these additions.

There are some caveats to doing this. Although you can overload a target already used by automake, it is often inadvisable, particularly in the topmost directory of a non-flat package. However, various useful targets have a ‘-local’ version you can specify in your ‘Makefile.in’. Automake will supplement the standard target with these user-supplied targets.

The targets that support a local version are all, info, dvi, check, install-data, install-

exec, and uninstall.

For instance, here is how to install a file in ‘/etc’:

     install-data-local:
             $(INSTALL_DATA) $(srcdir)/afile /etc/afile

Some targets also have a way to run another target, called a hook, after their work is done. The hook is named after the principal target, with ‘-hook’ appended. The targets allowing hooks are

install-data, install-exec, and dist.

For instance, here is how to create a hard link to an installed program:

     install-exec-hook:
             $(LN) $(bindir)/program $(bindir)/proglink

16 Distributing ‘Makefile.in’s

Automake places no restrictions on the distribution of the resulting ‘Makefile.in’s. We still encourage software authors to distribute their work under terms like those of the GPL, but doing so is not required to use Automake.

Some of the files that can be automatically installed via the ‘--add-missing’ switch do fall under the GPL; examine each file to see.

17 Some example packages

Here are some examples of how Automake can be used.

17.1 The simplest GNU program

hello is renowned for its classic simplicity and versatility. What better place to begin a tour? The below shows what could be used as the Hello distribution’s ‘Makefile.am’.

bin_PROGRAMS = hello
hello_SOURCES = hello.c version.c getopt.c getopt1.c getopt.h
hello_LDADD = @ALLOCA@
info_TEXINFOS = hello.texi
hello_TEXINFOS = gpl.texi
EXTRA_DIST = testdata
check-local: hello
        @echo expect no output from diff
        ./hello > test.out
        diff -c $(srcdir)/testdata test.out
        rm -f test.out

Of course, Automake also requires some minor changes to ‘configure.in’. The new ‘configure.in’ would read:

dnl Process this file with autoconf to produce a configure script.
AC_INIT(hello.c)
VERSION=1.3
AC_SUBST(VERSION)
PACKAGE=hello
AC_SUBST(PACKAGE)
AC_PROG_CC
AC_PROG_CPP
AC_PROG_INSTALL
AC_STDC_HEADERS
AC_HAVE_HEADERS(string.h fcntl.h sys/file.h)
AC_ALLOCA
AC_OUTPUT(Makefile)

If Hello were really going to use Automake, the ‘version.c’ file would probably be deleted, or changed so as to be automatically generated.

17.2 A tricker example

Here is another, trickier example. It shows how to generate two programs (ctags and etags) from the same source file (‘etags.c’). The difficult part is that each compilation of ‘etags.c’ requires different cpp flags.

bin_PROGRAMS = etags ctags
ctags_SOURCES =
ctags_LDADD = ctags.o

ctags_DEPENDENCIES = ctags.o
etags.o:
        $(COMPILE) -DETAGS_REGEXPS etags.c
ctags.o:
        $(COMPILE) -DCTAGS -o ctags.o etags.c

Note that ctags_SOURCES is defined to be empty – that way no implicit value is substituted. The implicit value, however, is used to generate etags from ‘etags.o’.

ctags_LDADD is used to get ‘ctags.o’ into the link line, while ctags_DEPENDENCIES exists to make sure that ‘ctags.o’ gets built in the first place.

This is a somewhat pathological example.

17.3 Automake uses itself

Automake, of course, uses itself to generate its ‘Makefile.in’. Since Automake is a shallow package, it has more than one ‘Makefile.am’. Here is the top-level ‘Makefile.am’:

## Process this file with automake to create Makefile.in
AUTOMAKE_OPTIONS = gnits
MAINT_CHARSET = latin1
PERL = @PERL@
SUBDIRS = tests
bin_SCRIPTS = automake
info_TEXINFOS = automake.texi
pkgdata_DATA = clean-kr.am clean.am compile-kr.am compile-vars.am \
compile.am data.am depend.am \
dist-vars.am footer.am header.am header-vars.am \
kr-vars.am libraries-vars.am \
libraries.am library.am mans-vars.am \
program.am programs.am remake-hdr.am \
remake-subd.am remake.am scripts.am subdirs.am tags.am tags-subd.am \
tags-clean.am \
texi-version.am texinfos-vars.am texinfos.am \
libraries-clean.am programs-clean.am data-clean.am \
COPYING INSTALL texinfo.tex \

ansi2knr.c ansi2knr.1 \
aclocal.m4
## These must all be executable when installed.
pkgdata_SCRIPTS = config.guess config.sub install-sh mdate-sh mkinstalldirs
CLEANFILES = automake
# The following requires a fixed version of the Emacs 19.30 etags.
ETAGS_ARGS = automake.in --lang=none \
 --regex='/^@node[ \t]+\([^,]+\)/\1/' automake.texi
## `test -x' is not portable.  So we use Perl instead.  If Perl
## doesn't exist, then this test is meaningless anyway.
# Check to make sure some installed files are executable.
installcheck-local:
$(PERL) -e "exit ! -x '$(pkgdatadir)/config.guess';"
$(PERL) -e "exit ! -x '$(pkgdatadir)/config.sub';"
$(PERL) -e "exit ! -x '$(pkgdatadir)/install-sh';"
$(PERL) -e "exit ! -x '$(pkgdatadir)/mdate-sh';"
$(PERL) -e "exit ! -x '$(pkgdatadir)/mkinstalldirs';"
# Some simple checks:
# * syntax check with perl4 and perl5.
# * make sure the scripts don't use 'true'
# * expect no instances of '${...}'
# These are only really guaranteed to work on my machine.
maintainer-check: automake check
$(PERL) -c -w automake
@if grep '^[^#].*true' $(srcdir)/[a-z]*.am; then \
  echo "can't use 'true' in GNU Makefile" 1>&2; \
  exit 1; \
else :; fi
@if test `fgrep '$${' $(srcdir)/[a-z]*.am | wc -l` -ne 0; then \
  echo "found too many uses of '\$${'" 1>&2; \
  exit 1; \
fi
if $(SHELL) -c 'perl4.036 -v' >/dev/null 2>&1; then \
  perl4.036 -c -w automake; \
else :; fi
# Tag before making distribution.  Also, don't make a distribution if
# checks fail.  Also, make sure the NEWS file is up-to-date.
cvs-dist: maintainer-check
@if sed 1q NEWS | grep -e "$(VERSION)" > /dev/null; then :; else \
  echo "NEWS not updated; not releasing" 1>&2; \
  exit 1; \
fi

     cvs tag `echo "Release-$(VERSION)" | sed 's/\./-/g'`
     $(MAKE) dist

As you can see, Automake defines many of its own rules, to make the maintainer’s job easier. For instance the cvs-dist rule automatically tags the current version in the CVS repository, and then makes a standard distribution.

Automake consists primarily of one program, automake, and a number of auxiliary scripts. Automake also installs a number of programs which are possibly installed via the ‘--add-missing’ option; these scripts are listed in the pkgdata_SCRIPTS variable.

Automake also has a ‘tests’ subdirectory, as indicated in the SUBDIRS variable above. Here is

‘tests/Makefile.am’:
     ## Process this file with automake to create Makefile.in
     AUTOMAKE_OPTIONS = gnits
     TESTS = mdate.test vtexi.test acoutput.test instexec.test checkall.test \
     acoutnoq.test acouttbs.test libobj.test proginst.test acoutqnl.test \
     confincl.test spelling.test prefix.test badprog.test depend.test
     EXTRA_DIST = defs

This is where all the tests are really run. ‘defs’ is an initialization file used by each test script; it is explicitly mentioned because automake has no way of automatically finding it.

17.4 A deep hierarchy

The GNU textutils are a collection of programs for manipulating text files. They are distributed as a deep package. The textutils have only recently been modified to use Automake; the examples come from a prerelease.

Here is the top-level ‘Makefile.am’:

     SUBDIRS = lib src doc man

In the ‘lib’ directory, a library is built which is used by each textutil. Here is ‘lib/Makefile.am’:

noinst_LIBRARIES = tu
EXTRA_DIST = rx.c regex.c
tu_SOURCES = error.h getline.h getopt.h linebuffer.h \
long-options.h md5.h regex.h rx.h xstrtod.h xstrtol.h xstrtoul.h \
error.c full-write.c getline.c getopt.c getopt1.c \
linebuffer.c long-options.c md5.c memchr.c safe-read.c \
xmalloc.c xstrtod.c xstrtol.c xstrtoul.c
tu_LIBADD = @REGEXOBJ@ @LIBOBJS@ @ALLOCA@

The ‘src’ directory contains the source for all the textutils – 23 programs in all. The ‘Makefile.am’ for this directory also includes some simple checking code, and constructs a ‘version.c’ file on the fly:

bin_PROGRAMS = cat cksum comm csplit cut expand fmt fold head join md5sum \
nl od paste pr sort split sum tac tail tr unexpand uniq wc
noinst_HEADERS = system.h version.h
DISTCLEANFILES = stamp-v version.c
INCLUDES = -I$(top_srcdir)/lib
LDADD = version.o ../lib/libtu.a
$(PROGRAMS): version.o ../lib/libtu.a
AUTOMAKE_OPTIONS = ansi2knr
version.c: stamp-v
stamp-v: Makefile
rm -f t-version.c
echo '#include <config.h>' > t-version.c
echo '#include "version.h"' >> t-version.c
echo 'const char *version_string = "'GNU @PACKAGE@ @VERSION@'";' \
>> t-version.c
if cmp -s version.c t-version.c; then \
  rm t-version.c; \
else \
  mv t-version.c version.c; \
fi
echo timestamp > $@
check: md5sum
./md5sum \
 --string="" \

Index of Targets 26

Index of Configure Variables and Macros

(Index is nonexistent)

Index of Targets

(Index is nonexistent)